Given the demand function P = 20 – 5Q, find the price elasticity of demand when price of the commodity is 5 Birr per unit. Mention if the demand is price elastic or inelastic at this point.
1
Expert's answer
2021-08-19T12:20:33-0400

P=20 - 5Q


5 = 20 - 5Q

5 - 20 = 5Q


-15 = 5Q


Q = 3



"E_d =\frac{dQ}{dP}.\frac{P}{Q}=-5×\frac{5}{3}=-8.333"


The demand price is inelastic